OpenStudy (anonymous):

Solve the equation by factoring 2x-17 = 9/x

ganeshie8 (ganeshie8):

\(\large 2x-17 = \frac{9}{x} \)

ganeshie8 (ganeshie8):

start by multiplying the x thru out the eq'n

ganeshie8 (ganeshie8):

\(\large x \times (2x-17 ) = x \times \frac{9}{x} \)

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

simplfy a bit, wat do u get

OpenStudy (anonymous):

Can you show me, I'm not sure

ganeshie8 (ganeshie8):

alright \(\large 2x^2-17x = 9 \) \(\large 2x^2-17x - 9 = 0 \)

ganeshie8 (ganeshie8):

u knw how to factor this ? :)

OpenStudy (anonymous):

I need my memory refreshed

OpenStudy (anonymous):

Hello?

ganeshie8 (ganeshie8):

wat grade u in ?

OpenStudy (anonymous):

I am taking College Algebra...I haven't used algebra in a long time

ganeshie8 (ganeshie8):

ohk

ganeshie8 (ganeshie8):

check this out http://www.purplemath.com/modules/solvquad.htm

ganeshie8 (ganeshie8):

u may ask me if u get any questions going thru that...

OpenStudy (anonymous):

Thanks for the link. I am still stuck on 2x^2 -17x-9 =0 Can you show me how to factor this part please?

ganeshie8 (ganeshie8):

u went thru the link ?

OpenStudy (anonymous):

I want to get a factor that equals 17 and - 9?

ganeshie8 (ganeshie8):

you're in right direction :)

ganeshie8 (ganeshie8):

basically we need to find two numbers :- 1) whose product equals 2x-9 = -18 2) whose sum equals -17

ganeshie8 (ganeshie8):

can u find them ?

ganeshie8 (ganeshie8):

observe that -18 and 1 will satisfy above

ganeshie8 (ganeshie8):

2x^2 -17x-9 =0 2x^2 -18x + 1x-9 =0

ganeshie8 (ganeshie8):

now factor first two terms, and last two terms

ganeshie8 (ganeshie8):

2x^2 -17x-9 =0 2x^2 -18x + 1x-9 =0 2x(x-9) + 1(x-9) = 0

ganeshie8 (ganeshie8):

notice that we can pull out (x-9)

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

2x^2 -17x-9 =0 2x^2 -18x + 1x-9 =0 2x(x-9) + 1(x-9) = 0 (x-9)(2x+1) = 0

ganeshie8 (ganeshie8):

next, use zero product property

ganeshie8 (ganeshie8):

2x^2 -17x-9 =0 2x^2 -18x + 1x-9 =0 2x(x-9) + 1(x-9) = 0 (x-9)(2x+1) = 0 x-9 = 0, 2x+1 = 0 x = 9 , x = -1/2

ganeshie8 (ganeshie8):

see if that makes more or less sense :)

OpenStudy (anonymous):

Yes, the last part I remember how to do. The factoring part is where I get confused at. Thanks so much for your help!

ganeshie8 (ganeshie8):

np :) just u have do few problems until u get confidence i guess... good luck !

OpenStudy (anonymous):

Thanks :)